# Mini-InternVL: A Flexible-Transfer Pocket Multimodal Model with 5% Parameters and 90% Performance

### v1

@@ 1 Introduction
In recent years, there have been significant advancements in multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6], which leverages the powerful capabilities of pre-trained large language models (LLMs) [7, 8, 9, 10, 11, 12] alongside vision foundation models (VFMs) [1, 13, 14]. These models undergo multi-stage training on extensive image-text data, which effectively aligns visual representations from VFMs with the latent space of LLMs, leading to promising performance in general vision-language understanding, reasoning, and interaction tasks. However, the large computational burden and the poor performance on long-tail domain-specific tasks hindered the widespread application of MLLMs in practical scenarios.
@@
The emergence of lightweight MLLMs  [15, 16, 17, 18] has provided a good balance between parameter size and performance, alleviating the reliance on expensive computing devices and fostering the development of various downstream applications. However, there are still several challenges: (1) Most existing MLLMs use visual encoders like CLIP [13], which are trained on Internet-domain image-text data and are aligned with BERT [19, 20]. As a result, these visual encoders are not capable of covering the extensive range of visual domains and are misaligned with LLMs’ representations. (2) To adapt MLLMs to specialized domains, existing methods mainly focus on modifying the model architectures, gathering extensive related training data, or customizing the training process for the target domain. There is still no consensus framework for LLMs’ downstream adaptation. Different domains have different model designs, data formats, and training schedules.
@@
To address these issues, there is a need for a strong visual encoder with comprehensive visual knowledge as well as a general transfer learning paradigm that allows for efficient application across downstream tasks in various domains at a low marginal cost.
@@
In this work, we introduce Mini-InternVL, a series of powerful pocket-sized MLLMs that can be easily transferred to various specialized domains. To this end, we first enhance the representational capabilities of a lightweight visual encoder. We initialize a 300M visual encoder using the weights from CLIP and apply knowledge distillation using InternViT-6B [1] as the teacher model. Subsequently, we develop Mini-InternVL series with 1 billion, 2 billion, and 4 billion parameters, by integrating the visual encoder with the pre-trained LLMs such as Qwen2-0.5B [11], InternLM2-1.8B [9], and Phi-3-mini [21], respectively. Benefiting from the robust visual encoder, Mini-InternVL exhibit excellent multimodal performance on general multimodal benchmarks like MMBench, ChartQA, and MathVista. Remarkably, compared with InternVL2-76B, the proposed Mini-InternVL-4B achieves 90% of the performance of larger counterparts while using to 5% fewer parameters, significantly reducing computational overhead.
@@
To further adapt our models to specific-domain downstream tasks, we introduce a straightforward yet effective transfer learning paradigm. Within this paradigm, we developed a unified transfer approach applicable to various downstream tasks, including autonomous driving, medical images, and remote sensing. This approach standardizes the model architecture, data format, and training schedule. The results demonstrate the effectiveness of this method in enhancing the model’s visual understanding and reasoning capabilities in domain-specific scenarios, enabling it to match the performance of proprietary commercial models within the target domains.
@@
(1) We propose Mini-InternVL, a powerful pocket multimodal model, that not only achieves robust multimodal performance with fewer than 4 billion parameters but also easily transfers to downstream tasks across various domains at low marginal cost.
(2) We develop several design features for Mini-InternVL, including a lightweight visual encoder—InternViT-300M, that is robust for various visual domains. Additionally, we introduce a simple but effective paradigm that standardizes model architecture, data format, and training schedule for effective downstream task transfer.
(3) We comprehensively evaluate our models through extensive experimentation on general and domain-specific benchmarks. The experimental results show that our multimodal models achieve 90% of the performance using significantly fewer parameters on general multimodal benchmarks. For specific domain tasks, with minimal computational cost for fine-tuning, they can rival closed-source commercial models. We conduct a series of ablation studies to explore the impact of data sample size on domain adaptation, hoping to provide insights into the application of MLLMs in specialized domains.
@@ 2 Related Works
Recently, some works, such as Fuyu [34],MoMa [35] and Chameleon [36], propose a visual encoder-free architecture. This type of architecture consists of a single Transformer model, which is used to process both visual and textual information simultaneously without requiring an additional encoder, making it more deployment-friendly. Despite these advancements, the heavy inference cost of these MLLMs hinders their application in downstream tasks. To address such issue, a series of ligthweight MLLMs, such as MiniCPM-V [15], are proposed. However, since most of them use CLIP-L [13] as visual encoder, which is only trained on natural image domain, these models are limited to the general domain and fail to generalize to other domain. In this work, we propose InternViT-300M, which is distilled from InternViT-6B and trained on a diverse image domain.
@@
From a vision-centric perspective, most MLLMs utilize vision models such as CLIP [13, 37] and SigLIP [14], which are trained on large-scale web image-text data. However, such visual encoders face significant limitations in terms of parameter scale and representational ability. Several studies have explored this issue. For instance, Tong et al. [38] identified significant differences in the visual patterns of CLIP and DINOv2 [39], leading to the development of a mixture-of-features module that integrates these two VFMs. LLaVA-HR [40] introduced a dual-branch vision encoder that employs CLIP-ViT for low-resolution pathways and CLIP-ConvNext for high-resolution pathways. Similarly, DeepSeek-VL [41] utilized a dual vision encoder design, incorporating SigLIP-L for low-resolution images and SAM-B [42] for high-resolution images.
However, these methods involve excessively complex pathways, which complicates the practical application of the models. Moreover, such approaches do not resolve the issue of visual encoders lacking comprehensive visual knowledge across various domains. In contrast, InternViT [1] implements progressive image-text alignment, and acquires representational capabilities across multiple domains by performing generative training on datasets spanning various fields. We propose injecting visual knowledge from the capable vision encoder into lightweight visual models, thus avoiding the computational expense associated with iterative generative pretraining.
@@
Figure 1: Training method and architecture of Mini-InternVL. Left: We employ InternViT-6B [1] as the teacher model to perform knowledge distillation on the student model. Right: Mini-InternVL adopts the ViT-MLP-LLM architecture similar to popular MLLMs [15, 41, 27, 28, 51, 52], combining InternViT-300M with a series of lightweight LLMs through an MLP projector. Here, we employ a simple pixel shuffle to reduce the number of visual tokens to one-quarter
@@ 3.1 Mini-InternVL
As shown in Figure 1, Mini-InternVL consists of three main components: InternViT, MLP Project, and LLMs. We employ InternViT-300M as our visual encoder, a lightweight vision model inherits the capabilities of a powerful vision encoder. Based on InternViT-300M, we develop three versions of Mini-InternVL: Mini-InternVL-1B, Mini-InternVL-2B, and Mini-InternVL-4B. Each version is respectively connected to the pre-trained Qwen2-0.5B [11], InternLM2-1.8B [9], and Phi-3-mini [21]. Similar to other open-source MLLMs [27, 2, 5, 52], Mini-InternVL employs an MLP projector to connect the visual encoder and the LLMs.
We adopt a dynamic resolution input strategy similar to that of InternVL 1.5 [2], which improves the model’s ability to capture fine-grained details. We also apply a pixel shuffle operation to reduce the number of visual tokens to one-quarter of the original. Consequently, in our model, a 448 448 image is represented by 256 visual tokens, enabling it to process up to 40 image crops (i.e., 4K resolution).
The training of Mini-InternVL consists of two stages: (1) Language-Image Alignment: We keep only the MLP component unfrozen during this stage. Following InternVL-1.5 [2], we use a diverse range of training datasets that encompass various tasks, including captioning, detection, grounding, and OCR. The diversity of these datasets ensures robust pre-training of Mini-InternVL, enabling the model to handle a variety of linguistic and visual elements across different tasks. (2) visual instruction tuning: We carefully select datasets to enhance the model s performance across a broad spectrum of multimodal tasks, similar to InternVL-1.5. These tasks include image captioning, chart interpretation, OCR, and cross-disciplinary reasoning. We conduct full-parameter fine-tuning with these datasets, further injecting world knowledge and teaching models to follow user instructions.
@@ 3.2 InternViT-300M
Table 1: Datasets used in knowledge distillation of visual encoder.
@@
Multidisciplinary CLEVR-Math/Super(en) [88, 89],GeoQA+ [90],
@@
Most existing MLLMs [15, 41, 27, 28, 51, 52] employ visual encoders that are trained on web-scale image-text paired data, such as CLIP, to obtain their representations. These encoders lack comprehensive knowledge of the visual world, which needs to be acquired through iterative generative pretraining in conjunction with LLMs. Unlike other approaches that enhance the visual foundation models by using auxiliary pathways [38, 41, 16], our method directly leverages a powerful vision model that has undergone generative training on diverse datasets to transfer knowledge to a lightweight vision model. Specifically, we use InternViT-6B as the teacher model and initialize the student model s weights using CLIP-ViT-L-336px. We align the representations of the student model with those of the teacher model by computing the negative cosine similarity loss between the hidden states of the last $K$ transformer layers. The resulting model is named InternViT-300M.
The primary goal of this knowledge transfer is to inherit the pretraining knowledge embedded in InternViT-6B. To achieve this, we curate a dataset sourced from a diverse range of publicly accessible resources, as detailed in Table 1. This dataset comprises four main types of data: natural images, OCR images, charts, and multi-disciplinary images. All images are resized to a resolution of 448 $\times$448, and dynamic resolution [2] is disabled for training efficiency. Ultimately, we develop a visual encoder, termed InternViT-300M, which is infused with diverse knowledge and is adaptable to various language models.
@@ 3.3 Domain Adaptation
Figure 2: The data format of our adaptation framework. We formulate other visual tasks (Image classification, region perception, multi-view images tasks, video related tasks and visual grounding) into VQA format in our framework.
@@
Instruction tuning is a crucial training stage to teach models to follow user instructions, of which the training data is formulated as visual question answering (VQA) and conversation format. VQA datasets of the downstream tasks, such as RSVQA [104] and PMC-VQA [95], are directly utilized as instruction-following data. For other conventional tasks, as shown in Figure 2, we formulate them into VQA format according to the following approaches separately:
@@
(3) Region Perception Tasks. Region-level conversation tasks are prevalent in specialized domains. These tasks involve supplying the model with spatial location information, in addition to the question input. The model is required to focus on objects within the specified attention region to generate a response. Specifically, there are two implementation methods. The first method involves directly annotating the location on the image using bounding boxes, masks, or contours, as illustrated in Region Perception2 of Figure 2. The second method denotes the object within the question by <box>[[x1,y1,x2,y2]]</box>, where the coordinates are normalized between 0 and 1000. This notation guides the model’s attention to specific regions within the image, enabling it to perform tasks such as region-level captioning and region-specific VQA.
For example, in the context of remote sensing, the objective is to enable the model to identify the object located within the coordinates [x1, y1, x2, y2]. To achieve this, we use a prompt such as “What object is in this location<box>[[x1, y1, x2, y2]]</box>” as the input instruction, with “<ref>object name</ref><box>[[x1, y1, x2, y2]]</box>” serving as the label.
@@
(4) Multi-View Images. In Autonomous driving, the images are captured from six different viewpoints. As shown in Figure 3, we effectively utilize dynamic resolution to accommodate this type of data. Specifically, InternVL supports splitting images into 448$\times$448-sized tiles based on their aspect ratio. Consequently, we resize each image to 896$\times$448 pixels and then, as illustrated, combine these images in a fixed sequence, resulting in a final resolution of 2688$\times$896. This means that the images are automatically processed into 12 tiles, and an additional thumbnail is added to provide the model with global context. Furthermore, we labeled each viewpoint image with text indicating its camera position, such as “CAM_FRON”.
(5) Video Frames. InternVL supports video frames in an interleaved image format. We represent the frame sequence using a template such as “Frame1: <img><IMG_CONTEXT></img> Frame2: <img><IMG_CONTEXT></img>.”, where <IMG_CONTEXT> denotes the image tokens. For each image at a resolution of 448$\times$448, the model can accommodate sequences of up to 40 frames.
@@
Table 2: Comparison with other models on multimodal benchmarks. We evaluate models using the InternVL and VLMEvalKit [106] repositories. AI2D [80], ChartQA [83], DocVQA [74], InfoVQA [82], and MMBench [107] are tested with InternVL, while MathVista [108] and OCRBench [109] use VLMEvalKit. For MMMU [110], we report scores from OpenCompass leaderboard. The Avg. Score is the average of the scores from all tested benchmarks, with the OCR-Bench score divided by 10. The values in parentheses represent the relative parameters and performance of Mini-InternVL compared to InternVL2-Llama3-76B [2], which is considered as 100%.

### v3

@@ 1 Introduction
In recent years, there have been significant advancements in multimodal large language models (MLLMs) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11], which leverages the powerful capabilities of pre-trained large language models (LLMs) [12, 13, 14, 15, 16, 17] alongside vision foundation models (VFMs) [1, 18, 19]. These models undergo multi-stage training on extensive image-text data, which effectively aligns visual representations from VFMs with the latent space of LLMs, leading to promising performance in general vision-language understanding, reasoning, and interaction tasks. However, the large computational burden and the poor performance on long-tail domain-specific tasks hinder the widespread application of MLLMs in practical scenarios.
@@
The emergence of lightweight MLLMs [20, 21, 22, 23, 24, 25] has provided a good balance between parameter size and performance, alleviating the reliance on expensive computing devices and fostering the development of various downstream applications. However, there are still several challenges: (1) Most existing MLLMs use vision encoders like CLIP [18], which are trained on Internet-domain image-text data and are aligned with BERT [26, 27]. As a result, these vision encoders are not capable of covering the extensive range of visual domains and are misaligned with LLMs’ representations. (2) To adapt MLLMs to specialized domains, existing methods mainly focus on modifying the model architectures, gathering extensive related training data, or customizing the training process for the target domain. There is still no consensus framework for LLMs’ downstream adaptation. Different domains have different model designs, data formats, and training schedules.
@@
To address these issues, there is a need for a strong vision encoder with comprehensive visual knowledge as well as a general transfer learning paradigm that allows for efficient application across downstream tasks in various domains at a low marginal cost.
@@
In this work, we introduce Mini-InternVL, a series of powerful pocket-sized MLLMs that can be easily transferred to various specialized domains. To this end, we first enhance the representational capabilities of a lightweight vision encoder. We initialize a 300M vision encoder using the weights from CLIP and apply knowledge distillation using InternViT-6B [1] as the teacher model. Subsequently, we develop Mini-InternVL series with 1 billion, 2 billion, and 4 billion parameters, by integrating the vision encoder with the pre-trained LLMs such as Qwen2-0.5B [16], InternLM2-1.8B [14], and Phi-3-Mini [28], respectively. Benefiting from the robust vision encoder, Mini-InternVL exhibits excellent multimodal performance on general multimodal benchmarks like MMBench [29], ChartQA [30], and MathVista [31]. Remarkably, compared with InternVL2-76B, the proposed Mini-InternVL-4B achieves 90% of the performance of larger counterparts while using only 5% of the parameters, significantly reducing computational overhead.
@@
To further adapt our models to specific-domain downstream tasks, we introduce a straightforward yet effective transfer learning paradigm. Within this paradigm, we develop a unified transfer approach applicable to various downstream tasks, including autonomous driving, medical images, and remote sensing. This approach standardizes the model architecture, data format, and training schedule. The results demonstrate the effectiveness of this method in enhancing the models’ visual understanding and reasoning capabilities in domain-specific scenarios, enabling them to match the performance of proprietary commercial models within the target domains.
@@
(1) We propose Mini-InternVL, a powerful pocket multimodal model, that not only achieves robust multimodal performance with only 4 billion parameters but also easily transfers to downstream tasks across various domains at low marginal cost.
(2) We develop several design features for Mini-InternVL, including a lightweight vision encoder—InternViT-300M, that is robust for various visual domains. Additionally, we introduce a simple but effective paradigm that standardizes model architecture, data format, and training schedule for effective downstream task transfer.
(3) We comprehensively evaluate our models through extensive experiments on general and domain-specific benchmarks. These results show that our multimodal models achieve 90% of the performance using significantly fewer parameters on general multimodal benchmarks. For specific domain tasks, with minimal computational cost for fine-tuning, they can rival closed-source commercial models. We conduct a series of ablation studies to explore the impact of data sample size on domain adaptation, hoping to provide insights into the application of MLLMs in specialized domains.
@@ 2 Related Works
Recently, some works, such as Fuyu [44], MoMa [45] and Chameleon [46], propose a vision encoder-free architecture. This type of architecture consists of a single transformer model, which is used to process both visual and textual information simultaneously without requiring an additional encoder, making it more deployment-friendly. Despite these advancements, the heavy inference cost of these MLLMs hinders their application in downstream tasks. To address such issue, a series of lightweight MLLMs, such as MiniCPM-V [20], are proposed. However, since most of them use CLIP-L [18] as the vision encoder, which is only trained on the natural image domain, these models are limited to the general domain and fail to generalize to other domains. In this work, we propose InternViT-300M, which is distilled from InternViT-6B and trained on a diverse set of image domains.
@@
From a vision-centric perspective, most MLLMs utilize vision models such as CLIP [18, 47] and SigLIP [19], which are trained on large-scale web image-text data. However, such vision encoders face significant limitations in terms of parameter scale and representational ability. Several studies have explored this issue. For instance, Tong et al. [48] identified significant differences in the visual patterns of CLIP and DINOv2 [49], leading to the development of a mixture-of-features module that integrates these two VFMs. LLaVA-HR [50] introduces a dual-branch vision encoder that employs CLIP-ViT for low-resolution pathways and CLIP-ConvNext for high-resolution pathways. Similarly, DeepSeek-VL [51] utilizes a dual vision encoder design, incorporating SigLIP-L for low-resolution images and SAM-B [52] for high-resolution images.
However, these methods involve excessively complex pathways, which complicates the practical application of the models. Moreover, such approaches do not resolve the issue of vision encoders lacking comprehensive visual knowledge across various domains. In contrast, InternViT [1] implements progressive image-text alignment, and acquires representational capabilities across multiple domains by performing generative training on datasets spanning various fields. We propose injecting visual knowledge from the capable vision encoder into lightweight visual models, thus avoiding the computational expense associated with iterative generative pre-training.
@@
Figure 1: Training method and architecture of Mini-InternVL. Left: We employ InternViT-6B [1] as the teacher model to perform knowledge distillation on the student model. Right: Mini-InternVL adopts the ViT-MLP-LLM architecture similar to popular MLLMs [20, 51, 37, 38, 61, 62], combining InternViT-300M with a series of lightweight LLMs through an MLP projector. Here, we employ a simple pixel unshuffle to reduce the number of visual tokens to one-quarter
@@ 3.1 Mini-InternVL
As shown in Figure 1, Mini-InternVL consists of three main components: visual encoder, MLP projector, and LLM. We employ InternViT-300M as our visual encoder, a lightweight vision model that inherits the capabilities of a powerful vision encoder. Based on InternViT-300M, we develop three versions of Mini-InternVL: Mini-InternVL-1B, Mini-InternVL-2B, and Mini-InternVL-4B, which are respectively connected to the pre-trained Qwen2-0.5B [16], InternLM2-1.8B [14], and Phi-3-mini [28]. Similar to other open-source MLLMs [37, 2, 5, 62], Mini-InternVL employs an MLP projector to connect the vision encoder and the LLMs.
We adopt a dynamic resolution input strategy similar to that of InternVL 1.5 [2], which improves the model’s ability to capture fine-grained details. We also apply a pixel unshuffle operation to reduce the number of visual tokens to one-quarter of the original. Consequently, in our model, a 448$\times$448 image is represented by 256 visual tokens, enabling it to process up to 40 image tiles (i.e., 4K resolution).
The training of Mini-InternVL consists of two stages: (1) Language-image alignment: We keep only the MLP component unfrozen during this stage. Following InternVL 1.5 [2], we use a diverse range of training datasets that encompass various tasks, including captioning, detection, grounding, and OCR. The diversity of these datasets ensures robust pre-training of Mini-InternVL, enabling the model to handle a variety of linguistic and visual elements across different tasks. (2) Visual instruction tuning: We carefully select datasets to enhance the model’s performance across a broad spectrum of multimodal tasks, similar to InternVL 1.5. These tasks include image captioning, chart interpretation, OCR, and cross-disciplinary reasoning. We conduct full-parameter fine-tuning with these datasets, further injecting world knowledge and teaching models to follow user instructions.
@@ 3.2 InternViT-300M
Table 1: Datasets used in knowledge distillation of vision encoder.
@@
Multidisciplinary CLEVR-Math/Super [97, 98], GeoQA+ [99],
@@
Most existing MLLMs [20, 51, 37, 38, 61, 62] employ vision encoders that are trained on web-scale image-text paired data, such as CLIP, to obtain their representations. These encoders lack comprehensive knowledge of the visual world, which needs to be acquired through iterative generative pre-training in conjunction with LLMs. Unlike other approaches that enhance the visual foundation models by using auxiliary pathways [48, 51, 21], our method directly leverages a powerful vision model that has undergone generative training on diverse datasets to transfer knowledge to a lightweight vision model. Specifically, we use InternViT-6B as the teacher model and initialize the student model’s weights using CLIP-ViT-L-336px. We align the representations of the student model with those of the teacher model by computing the negative cosine similarity loss between the hidden states of the last $K$ transformer layers. The resulting model is named InternViT-300M.
The primary goal of this knowledge transfer is to inherit the pre-training knowledge embedded in InternViT-6B. To achieve this, we curate a dataset sourced from a diverse range of publicly accessible resources, as detailed in Table 1. This dataset comprises four main types of data: natural images, OCR images, charts, and multi-disciplinary images. All images are resized to a resolution of 448 $\times$448, and dynamic resolution [2] is disabled for training efficiency. Ultimately, we develop a vision encoder, termed InternViT-300M, which is infused with diverse knowledge and is adaptable to various language models.
@@ 3.3 Domain Adaptation
Figure 2: The data format of our adaptation framework. We formulate other visual tasks (i.e., image classification, region perception, multi-view image tasks, video-related tasks, and visual grounding) into VQA format in our framework.
@@
Instruction tuning is a crucial training stage to teach models to follow user instructions, where the training data is formulated as visual question answering (VQA) and conversation format. VQA datasets of the downstream tasks, such as RSVQA [113] and PMC-VQA [104], are directly utilized as instruction-following data. For other conventional tasks, as shown in Figure 2, we formulate them into VQA format according to the following approaches separately:
@@
(3) Region Perception Tasks. Region-level conversation tasks are prevalent in specialized domains. These tasks involve supplying the model with spatial location information, in addition to the question input. The model is required to focus on objects within the specified attention region to generate a response. Specifically, there are two implementation methods. The first method involves directly annotating the location on the image using bounding boxes, masks, or contours, as illustrated in the “Region Perception2” of Figure 2. The second method denotes the object within the question by <box>[[x1,y1,x2,y2]]</box>, where the coordinates are normalized between 0 and 1000. This notation guides the model’s attention to specific regions within the image, enabling it to perform tasks such as region-level captioning and region-specific VQA.
For instance, in remote sensing applications, the goal is to train the model to identify objects within specific coordinates [x1, y1, x2, y2]. To achieve this, we use a prompt such as “What object is in this location<box>[[x1, y1, x2, y2]]</box>” as the input instruction, with “<ref>object name</ref><box>[[x1, y1, x2, y2]]</box>” serving as the label.
@@
(4) Multi-View Images. In autonomous driving, the images are captured from six different viewpoints. As shown in Figure 3, we effectively utilize dynamic resolution to accommodate this type of data. Specifically, Mini-InternVL supports splitting images into 448$\times$448-sized tiles based on their aspect ratio. Consequently, we resize each image to 896$\times$448 pixels and then, as illustrated, combine these images in a fixed sequence, resulting in a final resolution of 2688$\times$896. This means that the images are automatically processed into 12 tiles, and an additional thumbnail is added to provide the model with global context. Furthermore, we label each viewpoint image with text indicating its camera position, such as “CAM_FRON”.
(5) Video Frames. Mini-InternVL supports video frames in an interleaved image format. We represent the frame sequence using a template such as “Frame1: <img><IMG_CONTEXT></img> Frame2: <img><IMG_CONTEXT></img>.”, where <IMG_CONTEXT> denotes the image tokens. For each image at a resolution of 448$\times$448, the model can accommodate sequences of up to 40 frames.
@@
Table 2: Comparison with other models on multimodal benchmarks. We evaluate models using the InternVL and VLMEvalKit [115] repositories. AI2D [90], ChartQA [30], DocVQA [84], InfoVQA [92], and MMBench [29] are tested with InternVL, while MathVista [31] and OCRBench [116] use VLMEvalKit. For MMMU [117], we report scores from the OpenCompass leaderboard. The Avg. Score is the average of the scores from all tested benchmarks, with the OCRBench score divided by 10. The values in parentheses represent the relative parameters and performance of Mini-InternVL compared to InternVL2-Llama3-76B [2], which is considered as 100%.